OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by five.

OpenStudy (anonymous):

What does coterminal mean

OpenStudy (anonymous):

Isn't it when you add or subtract 360 degrees?

OpenStudy (anonymous):

no f7cking idea m8

Directrix (directrix):

Two I am thinking of are these: (1) pi/5 + 2 pi = pi/5 + 10 pi / 5 = (11 pi )/ 5 and (2) pi/5 - 2 pi = ?

OpenStudy (anonymous):

how did you get pi/5 + 10 pi/5?

Directrix (directrix):

10 pi/5 = 2 pi. Every 360 degrees (2 pi), the angles are coterminal. Consider an angle of 30 degrees on the unit circle. Add 360 degree and the angle 390 degrees will have a side terminal with the 30 degree angle. Read the information at this link: http://hotmath.com/hotmath_help/topics/coterminal-angles.html for an explanation.

OpenStudy (anonymous):

OHHHHH I get it! Thank you(:

Directrix (directrix):

You are welcome.